40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2619     Entry time: Fri Feb 19 16:40:43 2010
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: rearrange the optics on the end table 

Koji and kiwamu

The existing optics on the ETMX/ETMY end table were rearranged in this morning.

 


The main things we have done are -

1. relocation of the optical levers for ETMs ( as mentioned in koji's entry )

This relocation can make a space so that we can setup the green locking stuffs.

The optical path of the green locking is planed to start from the right top corner on the table, therefore we had to relocate the oplevs toward the center of the table.

2. relocation of the lens just before the tube

Because we are going to shoot the green beam into the arm cavity, we don't want to have any undesired lenses before the cavity.

For this reason we changed the position of the lens, it was standing just in front of the tube, now it's standing on the left side of the big mirror standing center top.

Since we did not find a significant change in its the spot size of the transmitted light, we did not change the position of all the TRANS_MON_PDs and its mirrors. And they are now well aligned.

Attachment1: ETMX end table

Attachment2: ETMY end table

Attachment 1: DSC_1202.JPG  2.647 MB  | Show | Hide all | Show all
Attachment 2: DSC_1207.JPG  2.696 MB  | Hide | Hide all | Show all
DSC_1207.JPG
ELOG V3.1.3-